# MathScape: Benchmarking Multimodal Large Language Models in Real-World Mathematical Contexts

### v5

@@ 1. Introduction
Large language models (LLMs) have demonstrated exceptional performance across diverse tasks spanning myriad domains (OpenAI, 2023a; Touvron et al., 2023). Based on LLMs, MLLMs (Zhao et al., 2023; Wu et al., 2023; Bai et al., 2024; An et al., 2025, 2024; Luo et al., 2024) also show strong understanding ability among different modalities (Liu et al., 2023b; Bai et al., 2023b). Among MLLMs, Vision-Language Models (VLMs) have demonstrated state-of-the-art performance in traditional multimodal tasks, including image classification (Chen et al., 2024), image understanding (Li et al., 2023a, b), and image captioning (Bai et al., 2023b). Furthermore, their strong language understanding capabilities enable high performance in text-rich tasks, such as visual question answering (Liu et al., 2023b, a) and image-text retrieval (Chen et al., 2024). Recently, MLLMs have also made significant strides in solving mathematical problems. Therefore, a well-designed benchmark is essential to systematically evaluate their mathematical reasoning abilities. Several benchmarks, such as MATH-V (Wang et al., 2024), MathVerse (Zhang et al., 2024), and MathVista (Lu et al., 2023), have been developed to assess the mathematical capabilities of MLLMs. While these works represent notable progress, existing benchmarks primarily rely on synthetic or digitally rendered mathematical problems. In contrast, real-world users often pose math questions based on photos of printed documents or screens, introducing additional challenges such as image quality variations and contextual ambiguity. To bridge this gap, it is essential to develop a comprehensive benchmark that better reflects real-world scenarios.
@@ 4.1. Experimental Setups
Models In our evaluation of multimodal LLMs, we focused on both open-source and closed-source models that rank among the top performers on major multimodal LLM leaderboards. We included 9 closed-source models, including GPT-4V (gpt-4-1106-preview), GPT-4-Turbo (gpt-4-turbo-2024-04-09) (OpenAI, 2023b), GPT-4o (gpt-4o-2024-08-06) (OpenAI, 2023b), GeminiPro (Reid et al., 2024), Claude-3-Opus, Baichuan-VL  (Yang et al., 2023), Qwen-Max (Bai et al., 2023a), Qwen-Plus  (Bai et al., 2023a), GLM4V (GLM et al., 2024). For open-source models, we evaluate 3 large models (¿20B parameters) such as Yi-VL-34B (Young et al., 2024), Qwen2-VL-Instruct-72B (Bai et al., 2023b), LLaVA-OneVision-72B (llava-onevision-qwen2-72b-ov-hf) (Li et al., 2024). Additionally, we evaluate 7 commonly used small open-source models, including Deepseek-VL2-4.5B (Lu et al., 2024), LLaVA-1.6-7B(Liu et al., 2024) Qwen2-VL-Instruct-7B (Bai et al., 2023b), LLaVA-OneVision-7B (llava-onevision-qwen2-7b-ov-hf) (Li et al., 2024) and Llama-3.2-11B-Vision (Touvron et al., 2023) and two math-specific MLLMs Math-LLaVA (Shi et al., 2024) and G-LLaVA-7B (Gao et al., 2023).
@@
Evaluation We evaluate all the aforementioned models on the MathScape benchmark. Additionally, we select two models—GPT-4o (gpt-4o-2024-08-06) (OpenAI, 2023b) and LLaVA-OneVision-72B (llava-onevision-qwen2-72b-ov-hf) (Li et al., 2024)—and assess their performance using a pure PDF file, without images captured by humans (which were intentionally included to increase task difficulty). These results highlight that strong performance on PDF-based inputs does not necessarily translate to human-captured images, underscoring the significance of our MathScape benchmark. Additional results on pure PDF inputs can be found in the Appendix. For the generated answers, we employ GPT-4V to segment and evaluate the responses produced by the selected models.

### v6

@@ 1. Introduction
Large language models (LLMs) have demonstrated exceptional performance across diverse tasks spanning myriad domains (OpenAI, 2023b; Touvron et al., 2023). Based on LLMs, MLLMs (Zhao et al., 2023; Wu et al., 2023; Bai et al., 2024; An et al., 2025, 2024; Luo et al., 2024) also show strong understanding ability among different modalities (Liu et al., 2023b; Bai et al., 2023b). Among MLLMs, Vision-Language Models (VLMs) have demonstrated state-of-the-art performance in traditional multimodal tasks, including image classification (Chen et al., 2024), image understanding (Li et al., 2023a, b), and image captioning (Bai et al., 2023b). Furthermore, their strong language understanding capabilities enable high performance in text-rich tasks, such as visual question answering (Liu et al., 2023b, a) and image-text retrieval (Chen et al., 2024). Recently, MLLMs have also made significant strides in solving mathematical problems. Therefore, a well-designed benchmark is essential to systematically evaluate their mathematical reasoning abilities. Several benchmarks, such as MATH-V (Wang et al., 2024), MathVerse (Zhang et al., 2024), and MathVista (Lu et al., 2023), have been developed to assess the mathematical capabilities of MLLMs. While these works represent notable progress, existing benchmarks primarily rely on synthetic or digitally rendered mathematical problems. In contrast, real-world users often pose math questions based on photos of printed documents or screens, introducing additional challenges such as image quality variations and contextual ambiguity. To bridge this gap, it is essential to develop a comprehensive benchmark that better reflects real-world scenarios.
@@ 4.1. Experimental Setups
Models In our evaluation of multimodal LLMs, we focused on both open-source and closed-source models that rank among the top performers on major multimodal LLM leaderboards. We included 9 closed-source models, including GPT-4V (gpt-4-1106-preview), GPT-4-Turbo (gpt-4-turbo-2024-04-09) (OpenAI, 2023a), GPT-4o (gpt-4o-2024-08-06) (OpenAI, 2023a), GeminiPro (Reid et al., 2024), Claude-3-Opus, Baichuan-VL  (Yang et al., 2023), Qwen-Max (Bai et al., 2023a), Qwen-Plus  (Bai et al., 2023a), GLM4V (GLM et al., 2024). For open-source models, we evaluate 3 large models (¿20B parameters) such as Yi-VL-34B (Young et al., 2024), Qwen2-VL-Instruct-72B (Bai et al., 2023b), LLaVA-OneVision-72B (llava-onevision-qwen2-72b-ov-hf) (Li et al., 2024). Additionally, we evaluate 7 commonly used small open-source models, including Deepseek-VL2-4.5B (Lu et al., 2024), LLaVA-1.6-7B(Liu et al., 2024) Qwen2-VL-Instruct-7B (Bai et al., 2023b), LLaVA-OneVision-7B (llava-onevision-qwen2-7b-ov-hf) (Li et al., 2024) and Llama-3.2-11B-Vision (Touvron et al., 2023) and two math-specific MLLMs Math-LLaVA (Shi et al., 2024) and G-LLaVA-7B (Gao et al., 2023).
@@
Evaluation We evaluate all the aforementioned models on the MathScape benchmark. Additionally, we select two models—GPT-4o (gpt-4o-2024-08-06) (OpenAI, 2023a) and LLaVA-OneVision-72B (llava-onevision-qwen2-72b-ov-hf) (Li et al., 2024)—and assess their performance using a pure PDF file, without images captured by humans (which were intentionally included to increase task difficulty). These results highlight that strong performance on PDF-based inputs does not necessarily translate to human-captured images, underscoring the significance of our MathScape benchmark. Additional results on pure PDF inputs can be found in the Appendix. For the generated answers, we employ GPT-4V to segment and evaluate the responses produced by the selected models.